# Quantum Search with the Signless Laplacian

### v1

@@ I Introduction
where the Hamiltonian of the signless Laplacian quantum walk is $H=-\gamma Q$. The signless Laplacian is used in spectral graph theory, such as to count the number of bipartite components [10]. We will show in the next section that the signless Laplacian quantum walk can arise in the Heisenberg (XYZ) model for a spin network with certain couplings between spins, such in layer antiferromagnetic materials. While the signless Laplacian quantum walk has been explored for state transfer [11, 12, 4, 13, 14], here we conduct the first algorithmic investigation of the signless Laplacian quantum walk. Namely, we explore quantum search [15] on the complete bipartite graph with multiple marked vertices, an example of which is shown in Fig. 1. The partite sets have $N_{1}$ and $N_{2}$ vertices, of which $k_{1}$ and $k_{2}$ are “marked” in the respective sets. The total number of vertices is $N=N_{1}+N_{2}$. The goal is to find a marked vertex by querying an oracle as few times as possible. In a continuous-time quantum walk, the oracle is an additional term the Hamiltonian, and so the task is to evolve the system for as little time as possible. Note the algorithmic component of this work is crucial—without the oracle, Laplacian and signless Laplacian quantum walks on bipartite graphs are identical in their probability distributions [4], but as we will show, the walks differ within the context of searching, i.e., when an oracle is present.
@@
Searching the complete bipartite graph with multiple marked vertices was explored for the Laplacian and adjacency quantum walks in [16], and our investigation shows that the signless Laplacian quantum walk searches differently from the other two. Table 1 summarizes the results that we will be deriving. In the first two rows, with the Laplacian quantum walk, the jumping rate $\gamma$ could be chosen so that the system entirely evolves from a uniform superposition over all the vertices ${\left|s\right\rangle}$ to only the marked vertices in one partite set, while another value of $\gamma$ causes it to entirely evolve to only the marked vertices in the other set [16]. In the next two rows, with the adjacency quantum walk, a single value of $\gamma$ causes the system to evolve from ${\left|s\right\rangle}$ to a mixture of marked and unmarked vertices in both partite sets with a success probability (probability of measuring the position of the walker at a marked vertex) of $1/2+\sqrt{N_{1}N_{2}}/N$; this same value of $\gamma$ causes the system to evolve from another state ${\left|s_{A}\right\rangle}$ (this is called ${\left|\sigma\right\rangle}$ in [16]) to a superposition over the marked vertices with certainty [16]. In this research, we show that the signless Laplacian exhibits a combination of the traits of the Laplacian and adjacency quantum walks, as summarized in the last four rows of the table. For some value of $\gamma$, the system partially evolves from the uniform state ${\left|s\right\rangle}$ to the marked vertices in one partite set, and it perfectly evolves a different initial state ${\left|s_{Q}\right\rangle}$ to the marked vertices. A different value of $\gamma$ partially and perfectly evolves these initial states to the marked vertices in the other partite set. All this assumes that $N_{1}\not\approx N_{2}$ so that the graph is not approximately regular. If $N_{1}\approx N_{2}$, then the graph is approximately regular, and the Laplacian, adjacency, and signless Laplacians all evolve identically (up to a global, ignorable phase), and so they all behave like the adjacency quantum walk in Table 1. Thus, we take $N_{1}\not\approx N_{2}$ for the rest of the paper, and from [16], this is quantified by $N_{1}$ and $N_{2}$ having a difference that scales at least as $\sqrt{N}$. The definitions of the three initial states, ${\left|s\right\rangle}$, ${\left|s_{A}\right\rangle}$, and ${\left|s_{Q}\right\rangle}$, are summarized in Table 2, and these will be described more fully in Sec. V.
@@ II.1 Adjacency Quantum Walk
Since $X_{i}X_{j}$ only acts on vertices $i$ and $j$, we can focus on the $Z$ basis states corresponding to vertices $i$ and $j$ and ignore the rest (i.e., trace them out). Then, $X_{i}X_{j}$ is equivalent to $X\otimes X$ acting on ${\left|0\right\rangle}_{z}{\left|0\right\rangle}_{z}$ if the excitation is neither at vertices $i$ nor $j$, ${\left|0\right\rangle}_{z}{\left|1\right\rangle}_{z}$ if the excitation is at vertex $j$, and ${\left|1\right\rangle}_{z}{\left|0\right\rangle}_{z}$ if the excitation is at vertex $i$. It cannot act on ${\left|1\right\rangle}_{z}{\left|1\right\rangle}_{z}$ since there is only one excitation. Working out each of the cases using (5),
@@
$(Y\otimes Y){\left|0\right\rangle}_{z}{\left|0\right\rangle}_{z}$ $=\left(i{\left|1\right\rangle}_{z}\right)\left(i{\left|0\right\rangle}_{z}\right)=-{\left|1\right\rangle}_{z}{\left|1\right\rangle}_{z},$
@@ III Search Algorithm
Starting with Fig. 3a, the solid black curve depicts $\gamma=0.001$, and the success probability stays near its initial value of $8/768\approx 0.01$. Increasing the jumping rate, when $\gamma=0.015$, the dashed red curve indicates that the success probability now rises a little. Continuing, when $\gamma=0.002$, the dotted green curve reveals that the success probability increases significantly, and measuring the location of the walker finds it at a marked vertex with high probability. Further still, when $\gamma=0.0025$, the dot-dashed blue curve shows that the success probability does not build up as well. When $\gamma=0.003$, the dot-dot-dashed orange curve reveals that the success probability is building up again. Continuing in Fig. 3b, when $\gamma=0.0035$, the solid black curve shows that the success probability is increasing still. The dashed red curve is $\gamma=0.004$, and the success probability has reached a high peak. Increasing $\gamma$ further to $0.0045$, $0.005$, and $0.0055$, the dotted green, dot-dashed blue, and dot-dot-dashed orange curves show that the success probability gets worse and worse. Altogether, these plots indicate that there are two values of $\gamma$, called critical values, for which the success probability is maximized.
@@ IV Analytical Proof
${\left|s\right\rangle}=\frac{1}{\sqrt{N}}\begin{pmatrix}\sqrt{k_{1}}\\ \sqrt{k_{2}}\\ \sqrt{N_{1}-k_{1}}\\ \sqrt{N_{2}-k_{2}}\\ \end{pmatrix},$ (13)
and the search Hamiltonian (10) is
$H=-\gamma\begin{pmatrix}\frac{1}{\gamma}+N_{2}&\sqrt{k_{1}k_{2}}&0&\sqrt{k_{1}N_{k2}}\\ \sqrt{k_{1}k_{2}}&\frac{1}{\gamma}+N_{1}&\sqrt{k_{2}N_{k1}}&0\\ 0&\sqrt{k_{2}N_{k1}}&N_{2}&\sqrt{N_{k1}N_{k2}}\\ \sqrt{k_{1}N_{k2}}&0&\sqrt{N_{k1}N_{k2}}&N_{1}\\ \end{pmatrix},$ (14)
where $N_{ki}=N_{i}-k_{i}$.
@@
${\left|a\right\rangle},$ $-1-N_{2}\gamma,$ (15)
${\left|b\right\rangle},$ $-1-N_{1}\gamma,$
${\left|u\right\rangle}=\frac{1}{\sqrt{N}}\left(\sqrt{N_{2}}{\left|c\right\rangle}+\sqrt{N_{1}}{\left|d\right\rangle}\right),$ $-N\gamma,$
@@ IV.1 Nondegenerate Case
${\left|\psi(t)\right\rangle}$ $\approx\frac{2\sqrt{N_{1}N_{2}}}{N}e^{i(1+N_{2}\gamma)t}{\left|u\right\rangle}$
$\quad+\frac{N_{2}-N_{1}}{N}e^{iN\gamma t}{\left|v\right\rangle}$
$=\frac{2\sqrt{N_{1}N_{2}}}{N}e^{i(1+N_{2}\gamma)t}\frac{1}{\sqrt{N}}\left(\sqrt{N_{2}}{\left|c\right\rangle}+\sqrt{N_{1}}{\left|d\right\rangle}\right)$
$\quad+\frac{N_{2}-N_{1}}{N}e^{iN\gamma t}\frac{1}{\sqrt{N}}\left(\sqrt{N_{1}}{\left|c\right\rangle}-\sqrt{N_{2}}{\left|d\right\rangle}\right).$
@@ IV.2 First Degenerate Case
$p_{a}(t)$ $=\frac{4N_{1}N_{2}}{N^{2}}\sin^{2}\left(\frac{\Delta E}{2}t\right),$
@@
As a check, note that $p_{a}(t)+p_{b}(t)+p_{c}(t)+p_{d}(t)=1$, since the total probability of the particle being at any vertex must be 1. We see that for large $N_{1}$ and $N_{2}$, there is asymptotically no probability of finding the particle at the marked $b$ vertices, but there is probability of finding it at the marked $a$ vertices, and this success probability reaches a maximum at time
@@
As a check, in Fig. 3a, the dot-dashed green curve corresponds to $\gamma=0.002$, which is roughly our derived critical value of $\gamma=1/N_{1}=1/512=0.00195$, and at roughly our derived runtime of $t_{*}=(\pi/2)\sqrt{N_{1}N/(k_{1}N_{2})}=(\pi/2)\sqrt{512\cdot 768/(3\cdot 256)}\approx 35.54$, the success probability roughly reaches our derived maximum of $p_{a}(t_{*})=4N_{1}N_{2}/N^{2}=4(512)(256)/768^{2}\approx 0.889$.
@@ V A Better Initial State and Search Algorithm
$Q=\begin{pmatrix}N_{2}&\sqrt{k_{1}k_{2}}&0&\sqrt{k_{1}(N_{2}-k_{2})}\\ \sqrt{k_{1}k_{2}}&N_{1}&\sqrt{k_{2}(N_{1}-k_{1})}&0\\ 0&\sqrt{k_{2}(N_{1}-k_{1})}&N_{2}&\sqrt{(N_{1}-k_{1})(N_{2}-k_{2})}\\ \sqrt{k_{1}(N_{2}-k_{2})}&0&\sqrt{(N_{1}-k_{1})(N_{2}-k_{2})}&N_{1}\\ \end{pmatrix},$
and the eigenvectors and eigenvalues of it are
${\left|v_{0}\right\rangle}=\frac{1}{\sqrt{N}}\left(\sqrt{k_{1}}{\left|a\right\rangle}-\sqrt{k_{2}}{\left|b\right\rangle}+\sqrt{N_{1}-k_{1}}{\left|c\right\rangle}-\sqrt{N_{2}-k_{2}}{\left|d\right\rangle}\right),$ $0,$
${\left|v_{1}\right\rangle}=\frac{1}{\sqrt{N_{2}}}\left(\sqrt{N_{2}-k_{2}}{\left|b\right\rangle}-\sqrt{k_{2}}{\left|d\right\rangle}\right),$ $N_{1},$
${\left|v_{2}\right\rangle}=\frac{1}{\sqrt{N_{1}}}\left(\sqrt{N_{1}-k_{1}}{\left|a\right\rangle}-\sqrt{k_{1}}{\left|c\right\rangle}\right),$ $N_{2},$
${\left|v_{3}\right\rangle}=\frac{1}{\sqrt{N}}\left(\sqrt{\frac{k_{1}N_{2}}{N_{1}}}{\left|a\right\rangle}+\sqrt{\frac{k_{2}N_{1}}{N_{2}}}{\left|b\right\rangle}+\sqrt{\frac{(N_{1}-k_{1})N_{2}}{N_{1}}}{\left|c\right\rangle}+\sqrt{\frac{(N_{2}-k_{2})N_{1}}{N_{2}}}{\left|d\right\rangle}\right),$ $N.$
@@
$p_{a}(t)$ $=\sin^{2}\left(\frac{\Delta E}{2}t\right),$

### v2

@@ I Introduction
where the Hamiltonian of the signless Laplacian quantum walk is $H=-\gamma Q$. The signless Laplacian is used in spectral graph theory, such as to count the number of bipartite components [10]. We will show in the next section that the signless Laplacian quantum walk can arise in the Heisenberg (XYZ) model for a spin network with certain couplings between spins, such in layer antiferromagnetic materials. While the signless Laplacian quantum walk has been explored for state transfer [11, 12, 4, 13, 14], here we conduct the first algorithmic investigation of the signless Laplacian quantum walk. Namely, we explore quantum search [15] on the complete bipartite graph with multiple marked vertices, an example of which is shown in Fig. 1. The partite sets have $N_{1}$ and $N_{2}$ vertices, of which $k_{1}$ and $k_{2}$ are “marked” in the respective sets. The total number of vertices is $N=N_{1}+N_{2}$. The goal is to find a marked vertex by querying an oracle as few times as possible. In a continuous-time quantum walk, the oracle is an additional term the Hamiltonian, and so the task is to evolve the system for as little time as possible. Note the algorithmic component of this work is crucial—without the oracle, Laplacian and signless Laplacian quantum walks on bipartite graphs are identical in their probability distributions [4], but as we will show, the walks differ within the context of searching, i.e., when an oracle is present. The complete bipartite graph was chosen because, due to the symmetry of the graph and the search algorithm, it is perhaps the easiest graph to analyze for which the three quantum walks differ.
@@
Searching the complete bipartite graph with multiple marked vertices was explored for the Laplacian and adjacency quantum walks in [16], and our investigation shows that the signless Laplacian quantum walk searches differently from the other two. Table 1 summarizes the results that we will be deriving. In the first two rows, with the Laplacian quantum walk, the jumping rate $\gamma$ could be chosen so that the system entirely evolves from a uniform superposition over all the vertices ${\left|s\right\rangle}$ to only the marked vertices in one partite set, while another value of $\gamma$ causes it to entirely evolve to only the marked vertices in the other set [16]. In the next two rows, with the adjacency quantum walk, a single value of $\gamma$ causes the system to evolve from ${\left|s\right\rangle}$ to a mixture of marked and unmarked vertices in both partite sets with a success probability (probability of measuring the position of the walker at a marked vertex) of $1/2+\sqrt{N_{1}N_{2}}/N$; this same value of $\gamma$ causes the system to evolve from another state ${\left|s_{A}\right\rangle}$ (this is called ${\left|\sigma\right\rangle}$ in [16]) to a superposition over the marked vertices with certainty [16]. In this research, we show that the signless Laplacian exhibits a combination of the traits of the Laplacian and adjacency quantum walks, as summarized in the last four rows of the table. For some value of $\gamma$, the system partially evolves from the uniform state ${\left|s\right\rangle}$ to the marked vertices in one partite set, and it perfectly evolves a different initial state ${\left|s_{Q}\right\rangle}$ to the marked vertices. A different value of $\gamma$ partially and perfectly evolves these initial states to the marked vertices in the other partite set. All this assumes that $N_{1}\not\approx N_{2}$ so that the graph is not approximately regular. If $N_{1}\approx N_{2}$, as quantified in [16] as $N_{1}$ and $N_{2}$ having a difference that scales less than $\sqrt{N}$, then the graph is approximately regular, and the Laplacian, adjacency, and signless Laplacians all evolve identically (up to a global, ignorable phase), and so they all behave like the adjacency quantum walk in Table 1. Thus, we take $N_{1}\not\approx N_{2}$ for the rest of the paper, and from [16], this is quantified by $N_{1}$ and $N_{2}$ having a difference that scales at least as $\sqrt{N}$. The definitions of the three initial states, ${\left|s\right\rangle}$, ${\left|s_{A}\right\rangle}$, and ${\left|s_{Q}\right\rangle}$, are summarized in Table 2, and these will be described more fully in Sec. V.
@@ II.1 Adjacency Quantum Walk
Since $X_{i}X_{j}$ only acts on vertices $i$ and $j$, we can focus on the $Z$ basis states corresponding to vertices $i$ and $j$ and ignore the rest. Then, $X_{i}X_{j}$ is equivalent to $X\otimes X$ acting on ${\left|0\right\rangle}_{z}{\left|0\right\rangle}_{z}$ if the excitation is neither at vertices $i$ nor $j$, ${\left|0\right\rangle}_{z}{\left|1\right\rangle}_{z}$ if the excitation is at vertex $j$, and ${\left|1\right\rangle}_{z}{\left|0\right\rangle}_{z}$ if the excitation is at vertex $i$. It cannot act on ${\left|1\right\rangle}_{z}{\left|1\right\rangle}_{z}$ since there is only one excitation. Working out each of the cases using (5),
@@
$(Y\otimes Y){\left|0\right\rangle}_{z}{\left|0\right\rangle}_{z}$ $=\left(i{\left|1\right\rangle}_{z}\right)\left(i{\left|1\right\rangle}_{z}\right)=-{\left|1\right\rangle}_{z}{\left|1\right\rangle}_{z},$
@@ III Search Algorithm
Starting with Fig. 3a, the solid black curve depicts $\gamma=0.001$, and the success probability stays near its initial value of $8/768\approx 0.01$. Increasing the jumping rate, when $\gamma=0.0015$, the dashed red curve indicates that the success probability now rises a little. Continuing, when $\gamma=0.002$, the dotted green curve reveals that the success probability increases significantly, and measuring the location of the walker finds it at a marked vertex with high probability. Further still, when $\gamma=0.0025$, the dot-dashed blue curve shows that the success probability does not build up as well. When $\gamma=0.003$, the dot-dot-dashed orange curve reveals that the success probability is building up again. Continuing in Fig. 3b, when $\gamma=0.0035$, the solid black curve shows that the success probability is increasing still. The dashed red curve is $\gamma=0.004$, and the success probability has reached a high peak. Increasing $\gamma$ further to $0.0045$, $0.005$, and $0.0055$, the dotted green, dot-dashed blue, and dot-dot-dashed orange curves show that the success probability gets worse and worse. Altogether, these plots indicate that there are two values of $\gamma$, called critical values, for which the success probability is maximized.
@@ IV Analytical Proof
${\left|s\right\rangle}=\frac{1}{\sqrt{N}}\begin{pmatrix}\sqrt{k_{1}}\\ \sqrt{k_{2}}\\ \sqrt{N_{1}-k_{1}}\\ \sqrt{N_{2}-k_{2}}\\ \end{pmatrix}.$ (13)
The adjacency matrix is
$A=\begin{pmatrix}0&\sqrt{k_{1}k_{2}}&0&\sqrt{k_{1}N_{k2}}\\ \sqrt{k_{1}k_{2}}&0&\sqrt{k_{2}N_{k1}}&0\\ 0&\sqrt{k_{2}N_{k1}}&0&\sqrt{N_{k1}N_{k2}}\\ \sqrt{k_{1}N_{k2}}&0&\sqrt{N_{k1}N_{k2}}&0\\ \end{pmatrix},$
where $N_{ki}=N_{i}-k_{i}$. For example, the first item of the last row, $\sqrt{k_{1}N_{k2}}=\sqrt{k_{1}(N_{2}-k_{2})}$ is obtained from the $k_{1}$ vertices of type $a$ multiplied by $\sqrt{N_{2}-k_{2}}/\sqrt{k_{1}}$ to convert between the normalizations of the ${\left|a\right\rangle}$ and ${\left|d\right\rangle}$ basis states. The degree matrix is $D=\text{diag}(N_{2},N_{1},N_{2},N_{1})$, and the Hamiltonian oracle is $-\sum_{i\text{ marked}}{\left|i\middle\rangle\middle\langle i\right|}=\text{diag}(-1,-1,0,0)$. Putting these together, the search Hamiltonian (10) is
$H=-\gamma\begin{pmatrix}\frac{1}{\gamma}+N_{2}&\sqrt{k_{1}k_{2}}&0&\sqrt{k_{1}N_{k2}}\\ \sqrt{k_{1}k_{2}}&\frac{1}{\gamma}+N_{1}&\sqrt{k_{2}N_{k1}}&0\\ 0&\sqrt{k_{2}N_{k1}}&N_{2}&\sqrt{N_{k1}N_{k2}}\\ \sqrt{k_{1}N_{k2}}&0&\sqrt{N_{k1}N_{k2}}&N_{1}\\ \end{pmatrix}.$ (14)
@@
${\left|a\right\rangle},$ $-1-\gamma N_{2},$ (15)
${\left|b\right\rangle},$ $-1-\gamma N_{1},$
${\left|u\right\rangle}=\frac{1}{\sqrt{N}}\left(\sqrt{N_{2}}{\left|c\right\rangle}+\sqrt{N_{1}}{\left|d\right\rangle}\right),$ $-\gamma N,$
@@ IV.1 Nondegenerate Case
${\left|\psi(t)\right\rangle}$ $\approx\frac{2\sqrt{N_{1}N_{2}}}{N}e^{i\gamma Nt}{\left|u\right\rangle}$
$\quad+\frac{N_{1}-N_{2}}{N}e^{0}{\left|v\right\rangle}$
$=\frac{2\sqrt{N_{1}N_{2}}}{N}e^{i\gamma Nt}\frac{1}{\sqrt{N}}\left(\sqrt{N_{2}}{\left|c\right\rangle}+\sqrt{N_{1}}{\left|d\right\rangle}\right)$
$\quad+\frac{N_{1}-N_{2}}{N}\frac{1}{\sqrt{N}}\left(\sqrt{N_{1}}{\left|c\right\rangle}-\sqrt{N_{2}}{\left|d\right\rangle}\right).$
@@ IV.2 First Degenerate Case
$p_{a}(t)$ $=\frac{4N_{1}N_{2}}{N^{2}}\sin^{2}\left(\frac{\Delta E}{2}t\right),$ (17)
@@
As a check, in Fig. 3a, the dot-dashed green curve corresponds to $\gamma=0.002$, which is roughly our derived critical value of $\gamma=1/N_{1}=1/512=0.00195$ We have reproduced this dot-dashed green curve in Fig. 5, and we have also added a plot of the asymptotic result $p_{a}(t)$ from (17) as a solid black curve, and we see strong agreement. As another check, note that $p_{a}(t)+p_{b}(t)+p_{c}(t)+p_{d}(t)=1$, since the total probability of the particle being at any vertex must be 1. We see that for large $N_{1}$ and $N_{2}$, there is asymptotically no probability of finding the particle at the marked $b$ vertices, but there is probability of finding it at the marked $a$ vertices, and this success probability reaches a maximum at time
@@
Figure 5: Success probability versus time for search on a complete bipartite graph of $N_{1}=512$ and $N_{2}=256$ vertices with $k_{1}=3$ and $k_{2}=5$ marked vertices, starting with the uniform state ${\left|s\right\rangle}$ and using a signless Laplacian quantum walk. The dotted green curve is $\gamma=0.002$ and is copied from Fig. 3a, and the solid black curve is the asymptotic formula $p_{a}(t)$ derived in (17).
As a check, in Fig. 5, the success probability reaches a maximum at our derived runtime of $t_{*}=(\pi/2)\sqrt{N_{1}N/(k_{1}N_{2})}=(\pi/2)\sqrt{512\cdot 768/(3\cdot 256)}\approx 35.54$, at which the success probability reaches our derived maximum of $p_{a}(t_{*})=4N_{1}N_{2}/N^{2}=4(512)(256)/768^{2}\approx 0.889$.
@@ V A Better Initial State and Search Algorithm
$Q=\begin{pmatrix}N_{2}&\sqrt{k_{1}k_{2}}&0&\sqrt{k_{1}N_{k2}}\\ \sqrt{k_{1}k_{2}}&N_{1}&\sqrt{k_{2}N_{k1}}&0\\ 0&\sqrt{k_{2}N_{k1}}&N_{2}&\sqrt{N_{k1}N_{k2}}\\ \sqrt{k_{1}N_{k2}}&0&\sqrt{N_{k1}N_{k2}}&N_{1}\\ \end{pmatrix},$
where $N_{ki}=N_{i}-k_{i}$ as before, and the eigenvectors and eigenvalues of $Q$ are
${\left|v_{0}\right\rangle}=\frac{1}{\sqrt{N}}\left(\sqrt{k_{1}}{\left|a\right\rangle}-\sqrt{k_{2}}{\left|b\right\rangle}\right.$
$\qquad\qquad\qquad\left.+\sqrt{N_{k1}}{\left|c\right\rangle}-\sqrt{N_{k2}}{\left|d\right\rangle}\right),$ $0,$
${\left|v_{1}\right\rangle}=\frac{1}{\sqrt{N_{2}}}\left(\sqrt{N_{k2}}{\left|b\right\rangle}-\sqrt{k_{2}}{\left|d\right\rangle}\right),$ $N_{1},$
${\left|v_{2}\right\rangle}=\frac{1}{\sqrt{N_{1}}}\left(\sqrt{N_{k1}}{\left|a\right\rangle}-\sqrt{k_{1}}{\left|c\right\rangle}\right),$ $N_{2},$
${\left|v_{3}\right\rangle}=\frac{1}{\sqrt{N}}\left(\sqrt{\frac{k_{1}N_{2}}{N_{1}}}{\left|a\right\rangle}+\sqrt{\frac{k_{2}N_{1}}{N_{2}}}{\left|b\right\rangle}\right.$
$\qquad\qquad\qquad\left.+\sqrt{\frac{N_{k1}N_{2}}{N_{1}}}{\left|c\right\rangle}+\sqrt{\frac{N_{k2}N_{1}}{N_{2}}}{\left|d\right\rangle}\right),$ $N.$
@@
$p_{a}(t)$ $=\sin^{2}\left(\frac{\Delta E}{2}t\right),$ (18)
@@
Figure 8: Success probability versus time for search on a complete bipartite graph of $N_{1}=512$ and $N_{2}=256$ vertices with $k_{1}=3$ and $k_{2}=5$ marked vertices, starting with ${\left|s_{Q}\right\rangle}$ and using a signless Laplacian quantum walk. The dotted green curve is $\gamma=0.002$ and is copied from Fig. 7a, and the solid black curve is the asymptotic formula $p_{a}(t)$ derived in (18).
Analogously to Fig. 5, as a check, in Fig. 8, we plot the dashed green curve from Fig. 7a along with the asymptotic formula $p_{a}(t)$ derived in (18) and see that there is strong agreement between the two.
@@
This results adds to a body of literature suggesting that starting in an eigenvector of the quantum walk operator (without the oracle) may be beneficial to search algorithms. This includes the result of Chakraborty et al. [22], which explored search on Erdös-Renyi random graphs using an adjacency quantum walk, and they showed that by starting in the principal eigenvector of the adjacency matrix, the success probability asymptotically reached $1$. In the setting of discrete-time, coined quantum walks, the abstract search algorithm [23] begins in an eigenvector of the quantum walk operator, although this success probability may not approach 1. Similarly, a discrete-time quantum walk searching the complete bipartite graph [24] evolves more smoothly from an eigenvector of the quantum walk operator than from the uniform state, but this success probability evolves around $1/2$ rather than $1$.